# LaSe-E2V: Towards Language-guided Semantic-Aware Event-to-Video Reconstruction

### v1

@@ Abstract
Event cameras harness advantages such as low latency, high temporal resolution, and high dynamic range (HDR), compared to standard cameras. Due to the distinct imaging paradigm shift, a dominant line of research focuses on event-to-video (E2V) reconstruction to bridge event-based and standard computer vision. However, this task remains challenging due to its inherently ill-posed nature: event cameras only detect the edge and motion information locally. Consequently, the reconstructed videos are often plagued by artifacts and regional blur, primarily caused by the ambiguous semantics of event data. In this paper, we find language naturally conveys abundant semantic information, rendering it stunningly superior in ensuring semantic consistency for E2V reconstruction. Accordingly, we propose a novel framework, called LaSe-E2V, that can achieve semantic-aware high-quality E2V reconstruction from a language-guided perspective, buttressed by the text-conditional diffusion models. However, due to diffusion models’ inherent diversity and randomness, it is hardly possible to directly apply them to achieve spatial and temporal consistency for E2V reconstruction. Thus, we first propose an Event-guided Spatiotemporal Attention (ESA) module to condition the event data to the denoising pipeline effectively. We then introduce an event-aware mask loss to ensure temporal coherence and a noise initialization strategy to enhance spatial consistency. Given the absence of event-text-video paired data, we aggregate existing E2V datasets and generate textual descriptions using the tagging models for training and evaluation. Extensive experiments on three datasets covering diverse challenging scenarios (e.g., fast motion, low light) demonstrate the superiority of our method. Dataset and code will be available upon acceptance. A demo video for the reconstructed results is attached in the supplementary material.
@@ 2 Related Works
Event-to-Video (E2V) Reconstruction. E2V reconstruction falls into two categories: model-based and learning-based methods. Model-based approaches [2, 33, 6, 41] exploit the correlation between events and intensity frames through hand-crafted regularization techniques. However, its reconstruction result is comparatively inferior to the more recent learning-based methods [38, 37, 42, 25, 11]. For example, E2VID [38, 37] used an Unet-like network with skip connections and ConvLSTM units to reconstruct videos from long event streams. Following E2VID, SPADE-E2VID [7] extended E2VID to enhance temporal coherence by feeding previously reconstructed frames into a SPADE block. HyperE2VID [11] introduced hypernetworks to generate per-pixel adaptive filters guided by a context fusion module. GANs, such as conditional GAN [48] and cycle-consistency GAN [54], are used to address this issue, but often exhibit blurry images with artifacts in non-activated areas. In summary, previous learning-based methods can only achieve plausible reconstructed results for this ill-posed problem because event data solely captures motion information without semantic context information. In this work, we explore the possibility of incorporating textual descriptions with semantic awareness for E2V reconstruction.
@@
Text-to-Video Diffusion Models. The success of Diffusion models [34, 40] in generating high-quality images from text prompts has advanced text-to-image (T2I) synthesis. Inspired by T2I synthesis, text-to-video (T2V) diffusion models, such as  [4, 14, 16, 17, 22, 29, 43, 47, 52] adapt T2I synthesis to video. Subsequent developments [17, 4, 13] focused on refining the temporal information interaction, such as optimizing temporal convolution or self-attention modules for motion learning. Meanwhile, one T2V research line focuses on enhancing controllability by incorporating additional conditions, thus addressing the text prompts’ ambiguity in motion, content, and spatial structure. For high-level video motion control, studies propose learning LoRA [18] layers for specific motion patterns [13, 59], or utilizing extracted trajectories [53], motion vectors [49], pose sequences [30] to guide the synthesis. For fine-grained spatial structure control, methods like Gen-1 [12], VideoComposer [49], and others [31, 55] leverage monocular depth, sketch, and image control models [31, 55] for flexible and controllable video generation [8, 13, 22, 58]. These methods achieve fine-grained controllability, focusing on flexible user instructions and diverse outcomes. In this study, we employ a basic T2V diffusion model with the 3D-UNet [47] architecture as the foundation for language-guided E2V reconstruction. Our framework prioritizes video fidelity by aligning motion details from event data and semantic insights from text, thus outperforming previous models, especially in extreme scenarios like fast motion and low light.
3 Preliminaries
Diffusion Model (DM) [44, 15]. Given a training sample $x_{0}\sim q(x_{0})$, DMs first obtain the noisy input through a forward diffusion process $q(x_{t}|x_{0},t),t\in\{1,2,...,T\}$ by using the parameterization trick [44] and gradually add Gaussian noise $\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ to the input $x_{t}=\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon$, where $\bar{\alpha}_{t}=\prod_{i=1}^{t}(1-\beta_{i})$. Note that $10<\beta_{1}<\beta_{2}<\ldots<\beta_{T}<$ 1 is a known variance schedule that controls the amount of noise added at each time step $t$. The diffusion model is then trained to approximate the backward process $p(x_{t-1}|x_{t})$ and recovers $x_{t-1}$ from $x_{t}$ using a denoising network $\epsilon_{\theta}(x_{t},\mathbf{c},t)$ by minimizing the mean squared error (MSE) between the predicted and target noise:
$\mathbb{E}_{x,\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I}),\mathbf{c},t}\left[\|\epsilon-\epsilon_{\theta}(x_{t},\mathbf{c},t)\|^{2}\right],$ (1)
which is denoted as $\epsilon$-prediction. Here, $\mathbf{c}$ denotes the (optional) conditional signal that DMs can be conditioned on. For our framework, $\mathbf{c}$ is an integration of the event data and text descriptions.
Latent Diffusion Model (LDM) [40]. Being variants of diffusion models, it uses a pre-trained encoder $\mathbf{E}_{I}$ to obtain a latent representation $z_{0}=\mathbf{E}_{I}(x_{0})$. LDMs then perform the forward process $q(z_{t}|z_{0},t)$ and the backward process $p_{\theta}(z_{t-1}|z_{t})$ in compressed latent space. The generated sample $\hat{x}_{0}$ can be obtained from the denoised latent using a pre-trained decoder $\hat{x}_{0}=\mathbf{D}_{I}(\hat{z}_{0})$.
@@ 4 The Proposed LaSe-E2V Framework
Event Representation. An event stream =$\mathcal{E}=\{e_{t_{k}}\}^{N_{e}}_{k=1}$, where $N_{e}$ represents the number of events. Each event $e_{t_{k}}\in\mathcal{E}$ is denoted as a four-element tuple $(x_{k},y_{k},t_{k},p_{k})$, indicating the spatial coordinates, timestamp and polarity respectively. To make the event stream compatible with the processing algorithms designed for frame-based vision, we convert an event stream $\mathcal{E}$ into the grid-like event voxel grid $V\in\mathbb{R}^{B\times H\times W}$ with $B$ time bins via temporal bilinear interpolation [61].
@@ 4.1 Overall Pipeline
Event Encoder $\mathbf{E}_{e}$. To adapt the event voxel grids to the latent space of $\mathbf{E}_{I}$, we first project it into a latent representation with an event encoder $\mathbf{E}_{e}(\mathbf{V})$. As illustrated in Fig. 2(a), we initially apply lightweight spatial 2D and temporal 3D convolution blocks to the input voxels $\mathbf{V}=\{V^{i}\}^{N}_{i=1}$. These blocks are designed to extract local spatial and temporal feature information, which is then processed through an attention block (AttBlock) for global temporal modeling. Subsequently, we concatenate the event latent representation with the noise latent representation along the channel dimension to serve as a rough condition for the model. Specially, we construct the input latent representation as $\bm{\hat{\epsilon}}=\{[\epsilon^{i},z_{e}^{i}]\}\in\mathbb{R}^{N\times(C^{\prime}+C_{e})\times H^{\prime}\times W^{\prime}}$, where $\epsilon^{i}$ and $z_{e}^{i}$ denote the noise latent and event feature corresponding to the $i$ frame with channel dimensions $C^{\prime}$ and $C_{e}$, respectively.
@@
Event-aware Mask Loss $l_{m}$. To optimize LaSe-E2V, we propose a novel event-aware mask loss $l_{m}$, in addition to the $\epsilon$-prediction loss (Eq. 1). The details of $l_{m}$ will be described in Sec. 4.3.
@@ 4.2 Event-guided Spatio-temporal Attention (ESA)
Event Spatial Attention. As illustrated in Fig. 2(b), the 2D spatial attention layer in the LDM U-Net contains a self-attention layer that operates on each frame independently and a cross-attention layer that operates between frames and the text embeddings. Given an intermediate hidden state $z^{i}$ of the $i^{th}$ frame, the self-attention operation is formulated as the attention between different spatial positions of $z^{i}$:
$\mathbf{Q}=\mathbf{W}^{Q}{z}^{i},\quad\mathbf{K}=\mathbf{W}^{K}{z}^{i},\quad\mathbf{V}=\mathbf{W}^{V}{z}^{i},$ (2)
$Attention(\mathbf{Q},\mathbf{K},\mathbf{V})=Softmax\left(\frac{\mathbf{Q}\mathbf{K}^{T}}{\sqrt{d}}\right)\mathbf{V},$ (3)
where $\mathbf{W}^{Q}$, $\mathbf{W}^{K}$ and $\mathbf{W}^{V}$ are learnable projection matrices, $d$ is the dimension of query and key.
Event data indicates abundant edge information in the spatial space, which is expected to directly constrain the structural information of the reconstructed video. To incorporate event information in the spatial attention module, we modify the key and value vectors in the self-attention layers to include features from the corresponding event latent $z_{e}^{i}$,
$\mathbf{Q}=\mathbf{W}^{Q}{z}^{i},\quad\mathbf{K}=\mathbf{W}^{K}[{z}^{i},{z}_{e}^{i}]_{s},\quad\mathbf{V}=\mathbf{W}^{V}[{z}^{i},{z}_{e}^{i}]_{s},$ (4)
where $[*]_{s}$ represents the concatenation operation on the spatial dimension. In this way, each spatial position in all frames gets access to the complete information from the corresponding event data, allowing fine-grained feature control in the spatial attention layers.
Event Temporal Attention. In addition to the spatial information, event data inherently represents the difference between the adjacent frames. As shown in Fig. 2(c), to effectively leverage the temporal constraint of the event data, we also augment the temporal self-attention layers to include the event feature. Formally, given an input hidden state $\mathbf{z}\in\mathbb{R}^{\in N\times C^{\prime}\times H^{\prime}\times W^{\prime}}$, where $N$ is the number of frames and $C^{\prime}$, $H^{\prime}$, $W^{\prime}$ correspond to the channel, height and width dimension of the hidden state tensor, we first reshape its height and width dimension to the batch dimension to form a new hidden state $\bar{\mathbf{z}}\in\mathbb{R}^{(H^{\prime}\times W^{\prime})\times N\times C^{\prime}}$. We then compute the query $\mathbf{Q}$, key $\mathbf{K}$ and value $\mathbf{V}$ from $z$ as:
$\mathbf{Q}=\mathbf{W}^{Q}{z}^{i},\quad\mathbf{K}=\mathbf{W}^{K}[{z}^{i},{z}_{e}^{i}]_{t},\quad\mathbf{V}=\mathbf{W}^{V}[{z}^{i},{z}_{e}^{i}]_{t},$ (5)
where $[*]_{t}$ represents the concatenation operation on the temporal dimension.
@@ 4.4 Event-aware Noise Initialization
During training, we construct the input latent by adding noise on the clean video latent. The noise schedule leaves some residual signal even at the terminal diffusion timestep $T$. As a result, the diffusion model has a domain gap to generalize the video during the inference time when we sample from random Gaussian noise without any real data signal. To solve this train-test discrepancy problem, during testing, we obtain the base noise by adding noise on event-accumulated frames using the forward process of DDPM [15]. The noise latent for frame $i$ can be expressed as:
@@ 5.1 Datasets and Implementation Details
Implementation Details Based on Stable Diffusion 2.1-base [40], we use a text-guided video diffusion model [39] to initialize our model, pre-trained on large-scale video datasets [1]. For each training video clip, we sample 16 frames and the corresponding event streams, with an interval of $1\leq v\leq 3$ frames. The input size is adapted to $256\times 256$. Following previous methods [51, 45], the data augmentation strategies include Gaussian noise, random flipping, and random pause. Our model is trained with the $\epsilon$-prediction objective overall U-Net parameters using a batch size of 3 and a learning rate of 5e-5 for 150k steps with 8 NVIDIA V100 GPUs. See more details in Appendix.

### v3

@@ Abstract
Event cameras harness advantages such as low latency, high temporal resolution, and high dynamic range (HDR), compared to standard cameras. Due to the distinct imaging paradigm shift, a dominant line of research focuses on event-to-video (E2V) reconstruction to bridge event-based and standard computer vision. However, this task remains challenging due to its inherently ill-posed nature: event cameras only detect the edge and motion information locally. Consequently, the reconstructed videos are often plagued by artifacts and regional blur, primarily caused by the ambiguous semantics of event data. In this paper, we find language naturally conveys abundant semantic information, rendering it stunningly superior in ensuring semantic consistency for E2V reconstruction. Accordingly, we propose a novel framework, called LaSe-E2V, that can achieve semantic-aware high-quality E2V reconstruction from a language-guided perspective, buttressed by the text-conditional diffusion models. However, due to diffusion models’ inherent diversity and randomness, it is hardly possible to directly apply them to achieve spatial and temporal consistency for E2V reconstruction. Thus, we first propose an Event-guided Spatiotemporal Attention (ESA) module to condition the event data to the denoising pipeline effectively. We then introduce an event-aware mask loss to ensure temporal coherence and a noise initialization strategy to enhance spatial consistency. Given the absence of event-text-video paired data, we aggregate existing E2V datasets and generate textual descriptions using the tagging models for training and evaluation. Extensive experiments on three datasets covering diverse challenging scenarios (e.g., fast motion, low light) demonstrate the superiority of our method.
@@ 2 Related Works
Event-to-Video (E2V) Reconstruction. E2V reconstruction falls into two categories: model-based and learning-based methods. Model-based approaches [2, 33, 6, 41] exploit the correlation between events and intensity frames through hand-crafted regularization techniques. However, its reconstruction result is comparatively inferior to the more recent learning-based methods [38, 37, 42, 25, 11]. For example, E2VID [38, 37] used an Unet-like network with skip connections and ConvLSTM units to reconstruct videos from long event streams. Following E2VID, SPADE-E2VID [7] extended E2VID to enhance temporal coherence by feeding previously reconstructed frames into a SPADE block. HyperE2VID [11] introduced hypernetworks to generate per-pixel adaptive filters guided by a context fusion module. GANs, such as conditional GAN [47] and cycle-consistency GAN [53], are used to address this issue, but often exhibit blurry images with artifacts in non-activated areas. In summary, previous learning-based methods can merely achieve plausible reconstructed results for this problem because event data only captures motion information without semantic context information. In this work, we explore the possibility of incorporating textual descriptions with semantic awareness for E2V reconstruction.
@@
Text-to-Video Diffusion Models. The success of Diffusion models [34, 40] in generating high-quality images from text prompts has advanced text-to-image (T2I) synthesis. Inspired by T2I synthesis, text-to-video (T2V) diffusion models, such as  [4, 14, 16, 17, 22, 29, 43, 46, 51] adapt T2I synthesis to video. Subsequent developments [17, 4, 13] focused on refining the temporal information interaction, such as optimizing temporal convolution or self-attention modules for motion learning. Meanwhile, one T2V research line focuses on enhancing controllability by incorporating additional conditions, thus addressing the text prompts’ ambiguity in motion, content, and spatial structure. For high-level video motion control, studies propose learning LoRA [18] layers for specific motion patterns [13, 58], or utilizing extracted trajectories [52], motion vectors [48], pose sequences [30] to guide the synthesis.
For fine-grained spatial structure control, methods like Gen-1 [12], VideoComposer [48], and others [31, 54] leverage monocular depth, sketch, and image control models [31, 54] for flexible and controllable video generation [8, 13, 22, 57]. Formally, given a video sample $x_{0}$, the latent diffusion model (LDM) [40] first encodes it into a latent feature $z_{0}=\mathbf{E}_{I}(x_{0})$. A noisy input is obtained based on the forward diffusion process by introducing Gaussian noise to the latent representation: $x_{t}=\sqrt{\bar{\alpha}_{t}}x_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon$, where $t=1,...,T$ and $T$ denotes the maximum timestep. $\bar{\alpha}_{t}=\prod_{i=1}^{t}(1-\beta_{i})$ is the coefficient that controls the noise strength. The diffusion model is trained by predicting the noise $\epsilon$ through a mean squared error:
$l_{\epsilon}=\|\epsilon-\epsilon_{\theta}(x_{t},\mathbf{c},t)\|^{2},$ (1)
where $\theta$ denotes the parameters of the U-Net of the diffusion model. $\mathbf{c}$ denotes the addition conditions (text, image, depth, et al.) to control the diffusion process.
These methods achieve fine-grained controllability, focusing on flexible user instructions and diverse outcomes. In this study, we employ a basic T2V diffusion model with the 3D-UNet [46] architecture as the foundation for language-guided E2V reconstruction. Our framework prioritizes video fidelity by aligning motion details from event data and semantic insights from text, thus outperforming previous models, especially in extreme scenarios like fast motion and low light.
@@ 3 The Proposed LaSe-E2V Framework
Event Representation. An event stream =$\mathcal{E}=\{e_{t_{k}}\}^{N_{e}}_{k=1}$ consists of $N_{e}$ events. Each event $e_{t_{k}}\in\mathcal{E}$ is represented as a tuple $(x_{k},y_{k},t_{k},p_{k})$, which denotes the pixel coordinates, timestamp, and polarity. To make the event stream compatible with the pre-trained diffusion model, we convert $\mathcal{E}$ into a grid-like event voxel grid $V\in\mathbb{R}^{B\times H\times W}$ with $B$ time bins using temporal bilinear interpolation [60].
@@ 3.1 Overall Pipeline
Event Encoder $\mathbf{E}_{e}$. To adapt the event voxel grids to the latent space of $\mathbf{E}_{I}$, we first project it into a latent representation with an event encoder $\mathbf{E}_{e}(\mathbf{V})$. As shown in Fig. 2(a), we initially apply lightweight spatial 2D and temporal 3D convolution blocks to the input voxels $\mathbf{V}=\{V^{i}\}^{N}_{i=1}$. These blocks are designed to extract local spatial and temporal feature information, which is then processed through an attention block (AttBlock) for global temporal modeling. Subsequently, we concatenate the event latent representation with the noise latent representation along the channel dimension to serve as a rough condition for the model. Specially, we construct the input latent representation as $\bm{\hat{\epsilon}}=\{[\epsilon^{i},z_{e}^{i}]\}\in\mathbb{R}^{N\times(C^{\prime}+C_{e})\times H^{\prime}\times W^{\prime}}$, where $\epsilon^{i}$ and $z_{e}^{i}$ denote the noise latent and event feature corresponding to the $i$ frame with channel dimensions $C^{\prime}$ and $C_{e}$, respectively.
@@
Event-aware Mask Loss $l_{m}$. To optimize LaSe-E2V, we propose a novel event-aware mask loss $l_{m}$, in addition to the $\epsilon$-prediction loss. The details of $l_{m}$ will be described in Sec. 3.3.
@@ 3.2 Event-guided Spatio-temporal Attention (ESA)
Event Spatial Attention. As illustrated in Fig. 2(b), the 2D spatial attention layer in the LDM U-Net integrates a self-attention mechanism that processes each frame individually and a cross-attention mechanism connecting frames with text embedding, which follows Stable Diffusion [40]. Intuitively, Event data indicates abundant edge information in the spatial space, which is expected to directly constrain the structural information of the reconstructed video. To incorporate event information in the spatial attention module, we concatenate the features from the event latent $z_{e}^{i}$ to the U-Net intermediate features $z$ to formulate a event-based spatial attention:
$z_{out}=Softmax\left(\frac{\mathbf{Q}\mathbf{K}_{e-s}^{T}}{\sqrt{d}}\right)\mathbf{V}_{e-s},$ (2)
where $\mathbf{Q}=\mathbf{W}^{Q}{z}^{i}$, $\mathbf{K}_{e-s}=\mathbf{W}^{K}[{z}^{i},{z}_{e}^{i}]_{s}$, $\mathbf{V}_{e-s}=\mathbf{W}^{V}[{z}^{i},{z}_{e}^{i}]_{s}$. $[*]_{s}$ represents the concatenation operation on the spatial dimension. This modification ensures that each spatial position in all frames accesses comprehensive information from the corresponding event data, enabling detailed structural feature control within the spatial attention layers.
Event Temporal Attention. In addition to the spatial information, event data inherently represents the difference between the adjacent frames, which serves as a strong constraint. As shown in Fig. 2(c), to effectively leverage the temporal constraint of the event data, we facilitate event features into temporal attention. Specially, given the intermediate features $z$, we first reshape the height and width dimensions into the batch dimension, forming a new hidden state $\bar{\mathbf{z}}\in\mathbb{R}^{(H\times W)\times N\times C}$. Then the event latent feature $z^{i}_{e}$ is employed to interact with the latent feature $\bar{z}$ in the temporal dimension through the temporal attention layer:
$z_{out}=Softmax\left(\frac{\mathbf{Q}\mathbf{K}_{e-t}^{T}}{\sqrt{d}}\right)\mathbf{V}_{e-t},$ (3)
where $\mathbf{Q}=\mathbf{W}^{Q}{\bar{z}}^{i}$, $\mathbf{K}_{e-t}=\mathbf{W}^{K}[{\bar{z}}^{i},{z}_{e}^{i}]_{t}$, $\mathbf{V}_{e-t}=\mathbf{W}^{V}[{\bar{z}}^{i},{z}_{e}^{i}]_{t}$. $[*]_{t}$ represents the concatenation operation on the temporal dimension.
@@ 3.4 Event-aware Noise Initialization
During training, we construct the input latent by adding noise on the clean video latent. The noise schedule leaves some residual signal even at the terminal diffusion timestep $T$. As a result, the diffusion model has a domain gap to generalize the video during the inference time when we sample from random Gaussian noise without any real data signal. To solve this train-test discrepancy problem, during test, we obtain the base noise by adding noise on event-accumulated frames using the forward process of DDPM [15]. The noise latent for frame $i$ can be expressed as:
@@ 4.1 Datasets and Implementation Details
Implementation Details Based on Stable Diffusion 2.1-base [40], we use a text-guided video diffusion model [39] to initialize our model, pre-trained on large-scale video datasets [1]. For each training video clip, we sample 16 frames and the corresponding event streams, with an interval of $1\leq v\leq 3$ frames. The input size is adapted to $256\times 256$. Following previous methods [50, 44], the data augmentation strategies include Gaussian noise, random flipping, and random pause. The model is trained using the $\epsilon$-prediction objective across all U-Net parameters, with a batch size of 3 and a learning rate of 5e-5 for 150k steps on 8 NVIDIA V800 GPUs.